# TextCenGen: Attention-Guided Text-Centric Background Adaptation for Text-to-Image Generation[Uncaptioned image]

### v4

@@ Introduction
Our goal is to maintain a smooth background in the text region denoted as $R$. As illustrated in Figure 2, during each time-step of the forward pass in the diffusion model, we modify the cross-attention maps at every layer. The cross-attention map is represented as $\mathcal{A}_{k}^{l}\in\mathbb{R}^{H\times W\times K}$, where $H\times W$ are the dimensions of $A^{l}$ at different scales, and $K$ signifies the maximum token length at layer $l$. The set $I$ consists of indices of tokens corresponding to areas outside the text in the prompt. We resize $R$ to align with the $HW$ dimensions. Subsequently, a new cross-attention map for each layer $l$ is computed as $\mathcal{A}_{k,\text{new}}^{l}=\{\lambda\times\mathcal{A}_{k}^{l}\odot(1-R)\,|\,\forall k\in O\}$, where $O$ represents the tokens needing editing. This procedure effectively redirects the model’s attention away from $R$, ensuring that the background in this region remains undisturbed and visually smooth. This spatially exclusive approach enhances the clarity and coherence of the generated images, particularly in areas designated for text insertion.
The final loss function of denoising process is defined as:
$\mathcal{L}=||\mathcal{A}_{ori}-\mathcal{A}_{res}||^{2}+\gamma\mathcal{L}_{norm}.$ (5)
In this case, $\mathcal{A}_{ori}$ and $\mathcal{A}_{res}$ stand for the original and altered versions of matrix $A$, correspondingly, while ||.||||.||| | . | | signifies the Euclidean distance among these cross-attention maps. Furthermore, the inclusion of the regularization term $\gamma\mathcal{L}_{norm}$ derived from self-guidance (Epstein et al. 2023), strives to prevent overfitting, where $\gamma$ denotes a weighting parameter.
@@
Dataset Metrics Stable Diffusion Dall-E 3 AnyText Desigen Ours
P2P Template Saliency IOU $\downarrow$ 29.89 52.64 54.34 30.62 28.37
TV Loss $\downarrow$ 14.11 18.02 22.55 13.70 13.29
VTCM $\uparrow$ 2.95 1.92 1.75 2.92 3.09
DiffuisonDB Saliency IOU $\downarrow$ 30.11 56.82 55.88 30.30 27.80
TV Loss $\downarrow$ 12.30 21.16 20.39 11.99 11.55
VTCM $\uparrow$ 3.19 1.67 1.79 3.20 3.37
Synthetics Prompt Saliency IOU $\downarrow$ 31.42 51.52 53.24 31.57 30.83
TV Loss $\downarrow$ 15.61 17.85 21.46 15.63 15.17
VTCM $\uparrow$ 2.73 1.98 1.79 2.67 2.77
Table 1: Quantitative comparison of our method with other methods (excluding CLIP Score). Bold indicates the best scores.
 Dataset Dall-E 3 AnyText Desigen Ours
P2P Template 2.53 0.32 0.55 0.30
DiffuisonDB 2.04 1.04 0.34 0.31
Synthetics Prompt 2.25 1.02 0.51 0.29
Experiments
@@
Our model is built with Diffusers. The pre-trained models are stable-diffusion-v1-5. While generating, the size of the output images is $512\times 512$. We use one A6000 and ten A40 GPUs for evaluation. Detailed parameter settings are provided in the appendix.
@@
Our evaluation contains 27,000 images generated from 2,700 unique prompts, each tested in ten different random region $R$. The dataset combined synthesized prompts generated by ChatGPT, and 700 prompts from the Prompt2Prompt template (Hertz et al. 2022) designed for attention guidance, focusing on specific objects and their spatial relationships. Additionally, we included 1,000 DiffusionDB prompts (Wang et al. 2022), chosen for their real-world complexity. This diverse and comprehensive dataset, spanning synthetic to user-generated prompts, provided a broad test ground to evaluate the efficacy of our TextCenGen method in various T2I scenarios.
Comparison with Existing Methods
We compared TextCenGen with several potential models to evaluate its efficiency. The baseline models included: Native Stable Diffusion (Rombach et al. 2022), Dall-E 3 (Ramesh et al. 2022), AnyText (Tuo et al. 2023) and Desigen (Weng et al. 2024). Dall-E used the prompt ”text-friendly in the position” to specify the region $R$. Similar to AnyText, we chose to randomly generate several masks in a fixed pattern across the map to simulate regions need to be edited. More detail can be found in appendix.
@@
To evaluate model performance, we used metrics assessing various aspects of the generated images. We proposed the CLIP Score Loss to evaluate the reduction in prompt semantic alignment for the training-free method compared to the original diffusion model. The CLIP Score (Hessel et al. 2021; Huang et al. 2021; Radford et al. 2021) measured semantic fidelity, ensuring images align with textual descriptions. The total variation (TV) loss (Rudin and Osher 1994; Jiang, Li, and Wang 2021) assessed the visual coherence and smoothness of the background in relation to text region $R$, crucial for harmonious compositions. Saliency Map Intersection Over Union (IOU) quantified focus and clarity around text areas. We propose the Visual-Textual Concordance Metric (VTCM) that combines a global metric increasing with value (CLIP Score) and local metrics that benefit from lower values (Saliency IOU and TV Loss) within $R$. The VTCM formula is:
$VTCM=\text{CLIP Score}\times(\frac{1}{\text{Saliency IOU}}+\frac{1}{\text{TV Loss}}).$ (6)
Our quantitative analysis in Table 1 and Table 2 shows that TextCenGen minimizes semantic loss while maintaining smoothness and saliency harmony within region R, even surpassing the latest Dalle-3. Scene text rendering and attention reduction in Desigen often focus on local attention, neglecting global semantics. Especially in Text in Object situation (see Figure 3), this weakens the ability to create whitespace. Our approach moves main objects away, then reduces attention to create space, resulting in more natural and harmonious text layouts with the highest VTCM.
@@
Following rising trend of multimodal large language model (MLLM) as judge methods (Chen et al. 2024; Wu et al. 2024), we present the MLLM-as-Judge ELO ranking for design appeal across different datasets in Table 3. The results demonstrate that the structured output from the GPT-4o with structured output provides consistent ratings, indicating that our method, along with Anytext and Dall-E, shows advantages in terms of design appeal. Interestingly, Dall-E excels in the LLM dataset that contains a higher proportion of natural landscapes. For detailed information on the prompts and evaluation methodology, please refer to the appendix.
Rank Method DDB P2P LLM
1 TextCenGen 702.21 752.85 122.96
@@
Our qualitative analysis, shown in Figure 4, involved a comparison across different models using prompts and positions within the same quadrant. Dall-E 3, which relies solely on text inputs, exhibited significant variability and could not consistently clear the necessary areas for text. Desigen reduces attention in region $R$, but this method is not always effective without pretrained graphic design-specific model weights, especially when region $R$ is within a main object, such as in the bicycle example. Introducing text to images was tricky, as TV Loss showed, but TextCenGen kept object detail and background quality well. It shows that our force-directed method effectively balances text and visuals in images.
User Study
To understand the importance of the text-friendliness issue and explore users’ subjective perceptions of the T2I Model’s results, we conducted a user study with 111 participants. We used the qualitative results (see Figure 4) to develop a questionnaire. Figure 5 illustrates that the prompt-image alignment and aesthetics of our results were well-received in human perception.
Ablation Study
@@
Our model’s force-directed feature is key to gently shifting where objects are placed. Without this, we might just bluntly edit the cross-attention map, which could mess up important parts of the picture. This part of our model helps us make sure we don’t ruin the image structure by harshly removing attention map from areas.
@@
Figure 5: User study of task importance and result evaluation. The left shows user perceptions of task encounter frequency (a1) and importance (a2), rating 5 as the highest. The right side details user rankings (b2) and the average times our results were chosen in a multiple-selection scenario (b1). The y-axis in b2 represents rankings from 1 to 5, demonstrating significant mean differences ($\alpha$=0.05) across three standards for our method.
w/o Both w/o FDG w/o SEC Ours
CLIP Score Loss $\downarrow$ - 2.2 1.71 0.32
TV Loss $\downarrow$ 14.39 12.44 12.76 13.67
Saliency IOU $\downarrow$ 30.32 28.56 28.61 28.34
VTCM $\uparrow$ 2.93 3.03 3.05 3.07
@@
Table 4: Ablation study result. We examine TextCenGen with or without the implementation of Force-Directed Cross-Attention Guidance (FDG) and Spatial Excluding Cross-Attention Constraint (SEC). The CLIP Score Loss of w/o both indicates the use of the original SD-1.5 without our training-free method, resulting in no loss (-).
Conclusion
@@
We present TextCenGen, a novel shift in T2I generation. This model abandons the traditional method of adapting text to pre-defined images for a more flexible, text-focused approach. TextCenGen modifies images to adapt text, employing force-directed cross-attention guidance to arrange whitespace. Furthermore, we have integrated a system to identify and relocate conflicting objects and a spatial exclusion cross-attention constraint for low saliency in whitespace areas.
@@
Our approach has certain limitations. The force-directed cross-attention guidance, assuming convexity and centering on object centroids, may not be suitable for non-convex shapes. It may lead to reduced size or fragmentation of objects. Future work will address these challenges to improve the quality of the output images. Additionally, we observed the generation of unintended objects in blank areas not mentioned in the original prompts.

### v5

@@ 3.2 Spatial Excluding Cross-Attention Constraint
Our goal is to maintain a smooth background in the text region denoted as $R$. As illustrated in Figure 2, during each time-step of the forward pass in the diffusion model, we modify the cross-attention maps at every layer. The cross-attention map is represented as $\mathcal{A}_{k}^{l}\in\mathbb{R}^{H\times W\times K}$, where $H\times W$ are the dimensions of $A^{l}$ at different scales, and $K$ signifies the maximum token length at layer $l$. The set $I$ consists of indices of tokens corresponding to areas outside the text in the prompt. We resize $R$ to align with the $H\times W$ dimensions. Subsequently, a new cross-attention map for each layer $l$ is computed as $\mathcal{A}_{k,\text{new}}^{l}=\{\mathcal{A}_{k}^{l}\odot(1-R)\,|\,\forall k\in O\}$, where $O$ represents the tokens needing editing. This procedure effectively redirects the model’s attention away from $R$, ensuring that the background in this region remains undisturbed and visually smooth. This spatially exclusive approach enhances the clarity and coherence of the generated images, particularly in areas designated for text insertion.
@@
4 Experiments
@@
Our model is built with Diffusers. The pre-trained models are stable-diffusion-v1-5 and stable-diffusion-v2-0. While generating, the size of the output images is $512\times 512$. We use one A6000 and ten A40 GPUs for evaluation. Detailed parameter settings are provided in the appendix.
@@
Our evaluation contains 27,000 images generated from 2,700 unique prompts, each tested in ten different random region $R$. The dataset combined synthesized prompts generated by ChatGPT, and 700 prompts from the Prompt2Prompt template (Hertz et al., 2022) designed for attention guidance, focusing on specific objects and their spatial relationships. Additionally, we included 1,000 DiffusionDB prompts (Wang et al., 2022), chosen for their real-world complexity. This diverse and comprehensive dataset, spanning synthetic to user-generated prompts, provided a broad test ground to evaluate the efficacy of our TextCenGen method in various T2I scenarios. Additionally, we constructed a targeted Desigen benchmark using 771 images from the Desigen dataset validation set (Weng et al., 2024), along with their corresponding static text masks, to evaluate performance on layout design-specific content. These 771 images were drawn from the original Desigen dataset of 53,577 usable images, where 52,806 were used for training the specialized Desigen model in Table 4.
4.2 Comparison with Existing Methods
We compared TextCenGen with several potential models to evaluate its efficiency. The baseline models included: Native Stable Diffusion (Rombach et al., 2022), Dall-E 3 (Ramesh et al., 2022), AnyText (Tuo et al., 2023) and Desigen (Weng et al., 2024). Dall-E used the prompt “text-friendly in the {position}" to specify the region $R$. Similar to AnyText, we chose to randomly generate several masks in a fixed pattern across the map to simulate regions need to be edited. More detail can be found in appendix.
@@ 4.2 Comparison with Existing Methods
To evaluate model performance, we used metrics assessing various aspects of the generated images. We proposed the CLIP Score Loss to evaluate the reduction in prompt semantic alignment for the training-free method compared to the vanilla diffusion model. The CLIP Score (Hessel et al., 2021; Huang et al., 2021; Radford et al., 2021) measured semantic fidelity, ensuring images align with textual descriptions. The total variation (TV) loss (Rudin & Osher, 1994; Jiang et al., 2021) assessed the visual coherence and smoothness of the background in relation to text region $R$, crucial for harmonious compositions. Saliency Map Intersection Over Union (IOU) (Qin et al., 2019) quantified the focus and clarity around text areas. We proposed the Visual-Textual Concordance Metric (VTCM), which combined a global metric increasing with value (CLIP Score) and local metrics that benefit from lower values (Saliency IOU and TV Loss) within $R$. The VTCM formula is:
VTCM=CLIP ScoreSaliency IOU+CLIP ScoreTV $\text{VTCM}=\frac{\text{CLIP Score}}{\text{Saliency IOU}}+\frac{\text{CLIP Score}}{\text{TV Loss}}$ (5)
Our quantitative analysis in Table 1 and Table 3 shows that TextCenGen minimizes semantic loss while maintaining smoothness and saliency harmony within region R, even surpassing the latest Dalle-3. Scene text rendering and attention reduction in Desigen often focus on local attention, neglecting global semantics. Especially in Text in Object situation (see Figure 3), this weakens the ability to create whitespace. Our approach moves main objects away, then reduces attention to create space, resulting in more natural and harmonious text layouts with the highest VTCM. The trade-off between background smoothness and semantic fidelity is shown in Figure 5.
We compared our method with the trained version of Desigen and retrieval-based methods like Text2Poster (Jin et al., 2022). Table 4 shows that the trained version of Desigen performs less effectively than our method despite using specialized graphic design data. As shown in Table 5, our plug-and-play method can be directly applied to pretrained model weights, yielding superior results compared to the same model using only attention reduction. When integrated with Desigen-Trained, our approach improves performance significantly, demonstrating that our training-free method can enhance specialized models without requiring additional training. As shown in Table 6, Text2Poster achieves competitive TV Loss on DiffusionDB, but our method provides better semantic fidelity (CLIP Score) and visual-textual coherence (VTCM). These results demonstrate the effectiveness of our training-free approach compared to both trained models and retrieval-based methods.
Figure 5: Performance trade-offs between different metrics. The dashed lines represent iso-utility curves, where points on the same curve indicate equivalent trade-off levels. Our method achieves a better balance between background smoothness and semantic fidelity. Higher utility curves (green) represent better overall performance.
@@
Following the rising trend of multimodal large language model (MLLM) as judge methods (Chen et al., 2024; Wu et al., 2024), we present the MLLM-as-Judge ELO ranking for design appeal across different datasets in Table 7. The results demonstrate that the structured output from the GPT-4o provides consistent ratings, indicating that our method, along with Anytext and Dall-E, shows advantages in terms of design appeal. Interestingly, Dall-E excels in the Synthesized Prompts dataset, which contains a higher proportion of natural landscapes. For detailed information on the prompts and evaluation, please refer to the appendix.
Rank Method DDB P2P SP
1 TextCenGen $\mathbf{702.21}$ $\mathbf{752.85}$ $\mathbf{122.96}$
@@
Our qualitative analysis, shown in Figure 4, involved a comparison across different models using prompts and positions within the same quadrant. Dall-E 3, which relied solely on text inputs, exhibited significant variability and could not consistently clear the necessary areas for text placement. Desigen reduced attention in region $R$, but this method was not always effective without pretrained graphic design-specific model weights, especially when region $R$ was within a main object, such as in the bicycle example. Introducing text to images was tricky, as TV Loss showed, but TextCenGen maintained object detail and background quality well. It showed that our force-directed method effectively balances text and visuals in images.
User Study.
To understand the importance of the text-friendliness issue and explore users’ subjective perceptions of the T2I Model’s results, we conducted a user study with 114 participants. We used the qualitative results (see Figure 4) to develop a questionnaire. Figure 6 illustrates that the prompt-image alignment and aesthetics of our results were well-received in human perception.
4.3 Ablation Study
@@
The force-directed module is key to gently shifting where objects are placed. Without this, we might just bluntly edit the cross-attention map, which could mess up important parts of the picture. This part of our model helps us make sure we don’t ruin the image structure by harshly removing attention map from areas.
@@
Figure 6: User study of task importance and result evaluation. The left shows user perceptions of task encounter frequency (a1) and importance (a2), rating 5 as the highest. The right side details user rankings (b2) and the average chosen times in a multiple-selection scenario (b1). The y-axis in b2 represents rankings from 1 to 5, demonstrating significant mean differences ($\alpha$=0.05) across three standards for all methods.
w/o All w/o FDG w/o SEC Ours
CLIPS Loss $\downarrow$ - 2.2 1.71 $\mathbf{0.32}$
TV Loss $\downarrow$ 14.39 12.44 12.76 $\mathbf{8.81}$
Saliency IOU $\downarrow$ 30.32 28.56 28.61 $\mathbf{22.86}$
VTCM $\uparrow$ 2.93 3.03 3.05 $\mathbf{4.4}$
@@
Table 8: Ablation study results. We examine TextCenGen with or without the implementation of Force-Directed Cross-Attention Guidance (FDG) and Spatial Excluding Cross-Attention Constraint (SEC). The CLIP Score Loss of w/o both indicates the use of the vanilla SD-1.5 without our training-free method, resulting in no loss (-).
5 Conclusion
@@
We present TextCenGen, a plug-and-play method for text-friendly image generation and requires no additional training while well balancing both semantic fidelity and visual quality. This method abandons the traditional method of adapting text to pre-defined images. TextCenGen modifies images to adapt text, employing force-directed cross-attention guidance to arrange whitespace. Furthermore, we have integrated a system to identify and relocate conflicting objects and a spatial exclusion cross-attention constraint for low saliency in whitespace areas.
@@
Our approach has certain limitations. The force-directed cross-attention guidance, which assumes convexity and centering on object centroids, may not be suitable for non-convex shapes. This may lead to reduced size or fragmentation of objects. Future work will address these challenges to improve the quality of the output images.
Acknowledgements
This work was supported by the NSSFC under Grant 22ZD05. We thank the anonymous reviewers and Dr. Sicheng Song for their valuable feedback and suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Task Introduction
Text-friendly images refer to images designed or selected with an emphasis on enhancing the readability and clarity of overlaid text. These images typically have simple, non-distracting backgrounds, a balanced color palette, and areas of negative space that can accommodate text without compromising visibility or design aesthetics. Common applications include marketing materials, presentations, and social media graphics where the text plays a crucial role in conveying the message. Key considerations for creating or selecting text-friendly images include contrast, alignment, and ensuring the image content does not compete with the overlaid text.
Task Name
Don’t Requires
Training
Don’t Requires
Annotation
Type of Layout
Specification
Required Number
of Anchors
Layout-to-Image (Rombach et al., 2022) X X Object $>5$
Text-Friendly Image Generation Space Region 1-2
Visual Text Generation Text 1-2
Table 9: Comparison of our task with existing tasks. Unlike layout-to-image tasks requiring training and intensive annotation, our method only needs space region annotation as input for downstream modifications. This makes it particularly suitable for applications such as dynamic wallpapers for mobile devices and e-commerce posters.
Appendix B Experiment Setting
Our proposed model is designed using the Diffusers library, specifically leveraging the stable-diffusion-v1-5 pre-trained models with DDPM Scheduler. The model generates images of dimensions $512\times 512$. In our method, we have set the force balance constant $\alpha$ to 0.5. The coefficients for regularization term $\gamma$ is fixed at 0.01. Within the detector, we upscale all cross-attention maps to a $64\times 64$ resolution. Additionally, we expand the height and width of the region $R$ by a margin of 0.06. During the first 20 steps, we identify conflicting objects when the Intersection over Union (IOU) exceeds 0.14. For the subsequent 30 steps, we initiate a push operation only if the average density inside region $R$ surpasses 0.8. Negative prompts are “monocolor, monotony, cartoon style, many texts, pure cloud, pure sea, extra texts, texts, monochrome, flattened, lowres, longbody, bad anatomy, bad hands, missing fingers, extra digit, fewer digits, cropped, worst quality, low quality”.
Our cross-attention replacement method requires less than 15GB of VRAM, making it feasible to run inference on consumer GPUs like the RTX 3090. For evaluation purposes, we utilized one NVIDIA A6000 and 8 H800 GPUs. The entire experimental assessment took approximately 96 hours to complete. Particularly on the A40 GPU, the image generation process, which includes both the original and the resultant images, took around 50 seconds for 50 steps of inference. In contrast, utilizing attention guidance with difftext results in a faster average inference time of approximately 30 seconds, as it only requires the inference of a single image.
B.1 Region Random Sampling Method
Our region random sampling method is a variation inspired by DiffText. It involves two predefined regions, measuring $160\times 64$ and $64\times 160$. During the evaluation, excluding Dall-E, we randomly select areas of these dimensions from the entire image. The output image of Dall-E is not $512\times 512$, so we proportionally scale down the corresponding regions to $116\times 46$ and $46\times 116$ while calculating metrics. Given that Dall-E operates exclusively on textual inputs as provided by the prompt, we intend to specify regions within Dall-E’s output at five distinct locations: left, right, bottom, top, and center.
B.2 Analysis of Text Box Shape Orientations
The shape orientation of text boxes was randomly generated following our region random sampling method, ensuring an unbiased experimental setup. To investigate potential biases related to text box orientations, we conducted additional analysis by separating results based on horizontal ($160\times 64$) and vertical ($64\times 160$) orientations. Table 10 shows the differences between horizontal and vertical orientations across different metrics and datasets, calculated as horizontal minus vertical values.
Dataset Metrics SD Dall-E 3 AnyText Desigen Ours
P2P Template Saliency IOU -1.54 0.17 -1.17 -0.57 0.31
TV Loss -4.83 5.89 -9.77 -1.26 3.77
VTCM -0.42 0.00 0.00 0.09 0.55
DiffusionDB Saliency IOU 0.83 -0.94 0.00 0.32 0.00
TV Loss 2.57 2.00 -3.48 2.49 -0.66
VTCM -0.32 1.76 -0.12 0.09 -1.14
Table 10: Differences between horizontal and vertical text box orientations across metrics and datasets. Positive values indicate better performance with horizontal boxes.
TextCenGen shows relatively consistent performance across orientations, with slightly better metrics for horizontal boxes in the P2P Template dataset. Different methods exhibit varying preferences for orientation, suggesting the underlying complexity of text-friendly image generation across different box shapes. The variance in performance across orientations is generally smaller for TextCenGen compared to baseline methods, indicating more robust handling of different text box shapes.
B.3 Evaluation Metrics
To evaluate model performance, we used 4 metrics assessing various aspects of the generated images. The CLIP Score, total variation (TV) loss, Saliency Map Intersection Over Union (IOU) and Visual-Textual Concordance Metric (VTCM).
  • CLIP Score is described as a metric used to measure the similarity between generated images and input prompts, employing off-the-shelf code referenced as (cli, 2022). Typically, the original image achieves the highest CLIP score when it’s associated with stable diffusion techniques. However, when blank areas are defined within an image, there might be a slight reduction in the CLIP Score. This score is crucial in determining how closely an AI-generated image aligns with the given textual prompt, playing a significant role in the overall evaluation of the image’s fidelity to the intended text description.
  • Total Variation Loss is a regularisation term commonly used in image processing tasks, particularly those involving image reconstruction or denoising. It is designed to encourage spatial smoothness in the output image while preserving important structural details. The total variation loss for a region $R$ can be computed as follows:
    $\text{TV}(R)=\sum_{i,j\in R}\sqrt{(\Delta_{x}R_{i,j})^{2}+(\Delta_{y}R_{i,j})^{2}}$
    Here, $\Delta_{x}R_{i,j}$ and $\Delta_{y}R_{i,j}$ represent the discrete differences in the horizontal (x-axis) and vertical (y-axis) directions, respectively, at pixel location $(i,j)$ within the region $R$. The term $(\Delta_{x}R_{i,j})^{2}+(\Delta_{y}R_{i,j})^{2}$ calculates the squared gradient magnitude at each pixel, and the summation is taken over all pixels within the region $R$.
    This formulation of the total variation loss ensures that the reconstructed or processed region $R$ does not have abrupt changes in pixel values, leading to a smoother and more visually appealing result.
  • Saliency Map Intersection Over Union is specifically tailored for assessing the overlap between a saliency map and a designated region within an image. The formula for this metric is given as:
    $\text{IoU}(S,R)=\frac{|S\cap R|}{|S\cup R|}$
    In this context, $S$ represents the saliency map, and $R$ denotes a specific region in the image. The term $|S\cap R|$ is the count of pixels that are common to both the saliency map $S$ and the region $R$, signifying their intersection. On the other hand, $|S\cup R|$ refers to the count of pixels present in either the saliency map $S$ or the region $R$, indicating their union.In our case, a lower IoU value is desirable as it indicates less overlap, aligning with our goal to ensure that the ROI is non-salient and distinct from the saliency map.
  • Visual-Textual Concordance Metric is formulated to assess the coherence between text prompts and generated images in AI-driven text-to-image synthesis. Defined as VTCM=CLIP Score×(1Saliency IOU+1TV $\text{VTCM}=\text{CLIP Score}\times\left(\frac{1}{\text{Saliency IOU}}+\frac{1}{\text{TV Loss}}\right)$, it combines three elements. The CLIP Score reflects the degree of match between the generated image and the text prompt, where higher scores indicate better alignment. The Saliency IOU (Intersection Over Union) in Region R measures how well the most salient parts of the image align with the specified region, with lower scores being better. The TV Loss in Region R assesses the smoothness or consistency in the image’s region, where lower TV Loss scores indicate a more uniform and less noisy region. The VTCM thus encourages the generation of images that are not only coherent with the text prompt but also exhibit focused quality in specified areas, aligning with the goal of creating text-friendly images.
Figure 8: The Limitation of Our Model.
Figure 9: The data generation method in Dall-E.
B.4 Details of Compared Methods
We compare the proposed model with Stable Diffusion, Dall-E, AnyText and DiffText. The details of compared methods are as follows:
  • Stable Diffusion, as detailed in Rombach et al. (2022) (Rombach et al., 2022), is an innovative open-source model. We utilize the widely accessible pre-trained model labeled as “runwayml/stable-diffusion-v1-5". For our experiments, we set the sampling steps to 50 and the classifier-free guidance scale at 7.5. The model serves both as the source for generating attention guidance from the origin image in our methodology and as one of the compared methods.
  • Dall-E (Ramesh et al., 2022) is a groundbreaking AI model developed by OpenAI, known for its capability to generate complex images from textual descriptions. Utilizing a variant of the GPT-3 architecture, Dall-E transforms text inputs into detailed and creative visual outputs, showcasing a deep understanding of both language and visual concepts. In our experiments, the output image of Dall-E is not $512\times 512$, so we proportionally scale down the corresponding regions to $116\times 46$ and $46\times 116$ while calculating metrics. To enhance the focus on text compatibility at these positions, we add the phrase ’text-friendly in the corresponding position’ directly into the original prompts, as depicted in Figure 9. Restricted to the purely textual input of Dall-E, the performance falters when the position falls between left and right and when the text lacks precision.
  • AnyText (Tuo et al., 2023) is a diffusion-based multilingual visual text generation and editing model that focuses on rendering accurate and coherent text in the image that outperformed all other approaches. We generate a fully black mask for that region and fully white for the rest of the image. Then we use this mask to replace the "draw_pos" in the input parameters, while keeping the other parameters unchanged.
  • Desigen (Weng et al., 2024) is an automatic template creation pipeline which generates background images as well as harmonious layout elements over the background as well as an iterative inference strategy to adjust the synthesized background and layout in multiple rounds is presented. For fair comparison, we adopted the following approaches: (1) We used random regions as layout elements, similar to other baselines. (2) Our comparison targets are training-free methods, so we used vanilla SD1.5 without any LoRA as the base weights. (3) Using Desigen’s attention reduction method on this base weight, we set the attention ratio within the region to 0.
B.5 MLLM-as-Judge ELO Ranking
To comprehensively assess the performance of TextCenGen against other baseline methods, we employed the Elo rating system, a method originally developed for ranking chess players but now widely used in various competitive contexts (Duan et al., 2024; Chen et al., 2024; Wu et al., 2024). In our evaluation framework, we apply a multi-modal large language model (MLLM) as a judge to compare TextCenGen with other baseline methods. The evaluation is based on the Elo rating system, commonly used in competitive ranking, which allows for continuous adjustments of scores as pairwise comparisons are made. Specifically, for each comparison between two methods, the MLLM assesses the "Design Appeal" of the generated outputs and determines a winner. The Elo method then updates the ratings of the two competing methods accordingly.
The Elo rating process involves calculating the expected scores $E_{A}$ and $E_{B}$ for methods $A$ and $B$, given their current ratings $R_{A}$ and $R_{B}$. The expected scores are computed using the formula:
$E_{A}=\dfrac{1}{1+10^{\dfrac{R_{B}-R_{A}}{400}}},\qquad E_{B}=\dfrac{1}{1+10^{\dfrac{R_{A}-R_{B}}{400}}}$ (6)
Based on the comparison outcome $S_{A}$ (where $S_{A}=1$ if method $A$ wins, or $S_{A}=0$ if it loses), the ratings are updated using:
$R_{A}^{\prime}$ $=R_{A}+K\times(S_{A}-E_{A})$
$R_{B}^{\prime}$ $=R_{B}+K\times((1-S_{A})-E_{B})$
Here, $K$ represents the adjustment factor, set to 32 in our experiments, which determines the sensitivity of rating changes.
In practice, our evaluation system iterates through each dataset, initializing the Elo scores for all methods and adjusting them dynamically as new comparison results are processed. After all comparisons, the methods are ranked based on their aggregated Elo scores, providing insights into the relative strengths of TextCenGen and other baselines in terms of "Design Appeal." This approach ensures a consistent and scalable evaluation across diverse datasets while reflecting the preferences derived from the MLLM’s judgments.
Figure 10: More results of elo mllm judge.
Figure 11: More results of the proposed method. The first and third lines display the original images, while the second and fourth lines exhibit the result images.
Appendix C Influence of the force balance constant
$\alpha$ ensures that the forces exerted do not exceed a practical threshold, thereby maintaining visual equilibrium in complex scenes (shown in Figure 12).
Figure 12: Influence of the force balance constant. The graph illustrates the effects of four different values for $\alpha$ on $F_{rep}$ versus the final force. The curves reveal that as $\alpha$ increases, the rate of convergence to the limit progresses more rapidly.
Appendix D More Results of Proposed Method
We present additional outcomes of the proposed methodology. Referring to Figure 11, our model produces 12 instances drawn from three distinct datasets. The figure also illustrates the strong performance of our approach across diverse scenarios. These encompass urban landscapes, natural scenes, still life, Valentine’s Day greeting cards, and environments from both 3D and 2D video games.
D.1 More Result of Ablation Study
Figure 13 presents the results of our ablation study. This study was designed to elucidate the individual contributions of the various components embedded within our proposed model. As part of our comprehensive analysis, we evaluated the three main contributors: (1) the impact of the force-directed component, (2) the effectiveness of the Spatial Excluding Cross-Attention Constraint.
Figure 13: The results of ablation. We examine TextCenGen with or without the implementation of Force-Directed Cross-Attention Guidance (FDG) and Spatial Excluding Cross-Attention Constraint (SEC). The red-dotted area denotes the target area preserved for either text or icon images. Images produced with both FDG and SEC yield outcomes identical to those created using TextCenGen alone. Conversely, images created without FDG and SEC are comparable to those derived from the original stable diffusion model.
D.2 Compatible with Lora Checkpoint
Figure 1 demonstrates the effect of using LoRA by utilizing the LoRA rev-animated model from Civitai for an animated style, while the qualitative evaluation uses original SD weights for comparison.
Appendix E Limitations of Our Model
Our approach could result in unexpected changes, exemplified by the left of Figure 8, where one might initially think the figure shifts downward, rather than the forehead area. We noted the emergence of unintended objects within the empty spaces on the right of Figure 8, spaces which the original prompts did not specify.